# Towards Internet-Scale Training For Agents

### v1

@@ Appendix G Agent & Judge System Prompts
We provide the system prompt used with our agent below. This prompt is released in our official code, alongside the observation processor that maps webpage DOM to a compact markdown format, referenced in the system prompt.
You are a helpful assistant operating my web browser. I will show you webpages formatted in markdown, and I want your help to complete a web navigation task. Read the webpage, and respond with an action in JSON to interact with the page, and help me complete the task.
## Formatting The Response
Respond with actions in the following JSON schema:
@@
"target_element_id": int
Here is what each key means:
- action_key‘: The action to perform.
- action_kwargs‘: Named arguments for the action.
- target_element_id‘: The id of the element to perform the action on.
## Available Actions
Im using playwright, a browser automation library, to interact with the page. Im parsing the value assigned to action_key into a method call on the page object, or an element object specified by the value assigned to target_element_id‘. Here are the available actions:
@@
Suppose you want to click the link ‘[id: 5] Sales link‘:
@@
Suppose you want to hover over the image ‘[id: 2] Company Logo image‘:
@@
### Example Fill Action
Suppose you want to fill the input ‘[id: 13] "Name..." (Enter your name text field)‘ with the text John Doe‘:
@@
Suppose you want to select the option red from the dropdown ‘[id: 67] "blue" (select a color dropdown)‘:
@@
Suppose you want to open google search:
@@
- stop‘: Stop the browser when the task is complete, or the answer is known.
- answer‘: Optional answer if I requested one.
@@
"answer": "Im done!"
@@
Thanks for helping me perform tasks on the web, please follow the instructions carefully. Start your response with an explanation in 50 words, and choose exactly one action you would like to perform.
We also provide the system prompt used with out LLM judge. The system prompt instructs the judge to predict a json-formatted dictionary that contains a “success” key, and an “on_right_track” that represent the estimated probability that the task is successful, and that the agent is on the right track towards solving the task, respectively. These distinctions are adapted from Koh et al. (2024b), and help us filter for high-quality training data by distinguishing trajectories that were solved by the agent’s own actions from trajectories that were solved by chance.
You are a helpful assistant providing feedback on a web automation script. I will show you a list of previous actions, the current webpage formatted in markdown, and the proposed next action. I want your help evaluating the proposed action, to determine if the desired task is complete, or if we are on the right track towards future completion.
## Reading The Action Schema
You will see actions in the following JSON schema:
@@
Here is what each key means:
- action_key‘: The action to perform.
- action_kwargs‘: Dictionary of arguments for action.
- target_element_id‘: The id of the element to perform the action on.
## Available Actions
Im using playwright, a browser automation library, to interact with the page. Im parsing the value assigned to action_key into a method call on the page object, or an element specified by the value assigned to target_element_id‘. Here is an example action:
@@
Here is an example where the script clicked the link ‘[id: 5] Sales link‘:
@@
Here is an example where the script selected the option red from the dropdown ‘[id: 67] "blue" (select a color dropdown)‘:
@@
Here is an example where the script opened google search:
@@
Here is an example where the script stopped with the message "Im done!":
@@
"answer": "Im done!"
@@
## Formatting The Response
Think step by step, and start your response with an explanation of your reasoning in 50 words. Then, provide an evaluation in the following JSON schema:
@@
"on_right_track": float,
Here is what each key means:
- success‘: What is the probability the desired task has been completed successfully, rated from 0.0 (not possible) to 1.0 (absolutely certain)?
- on_right_track‘: What is the probability the script is on the right track towards a future success, rated from 0.0 (not possible) to 1.0 (absolutely certain)?
Thanks for helping me evaluate the script, please follow the instructions carefully. Start your response with a step by step explanation. Then, provide an evaluation in the JSON schema above.
@@ Appendix H Details For Training Agents
To understanding the utility of the generated data for training agents, we filter the data, and compare our filtered data to human demonstrations on the WebLINX (Lù et al., 2024) and Mind2Web (Deng et al., 2023) benchmarks. In particular, we sweep over different sizes of random subsets of human actions, from 32 to 256, which helps us understand the value of synthetic data generated from the InSTA pipeline versus different scales of human data. We then fine-tune models based on google/flan-t5-large for Mind2Web, and meta-llama/Llama-3.1-8B-Instruct for WebLINX using official fine-tuning code released with corresponding benchmarks. We employ identical training hyperparameters to those used in Lù et al. (2024) and Deng et al. (2023) to ensure that our results are comparable to previous work. Section H reports performance on the official test_web split of the WebLINX benchmark, and the official test_website split of the Mind2Web benchmark, where agents are tested on previously unobserved websites.
In order to prepare our data, we employ three filtering rules. In the first rule, we filter for data where the agent was predicted to have succeeded at the task with conf = 1, and was predicted to be on the right track with conf = 1. This filtering rule is motivated by our findings in Section 5, where we found that our LLM judge based on Llama 3.1 70B has an accuracy up to 93.1% at detecting successful trajectories for its predictions with conf = 1. Filtering based on both “success” and “on_right_track” conditions is essential to obtain data where the agent directly caused the task to be solved, rather than the task being solved by external conditions. The next filtering rule we use is to select trajectories with at least three actions, which helps create training data that is not too easy (i.e. not solved after just one or two actions). Finally, we select tasks where the agent did not encounter any errors during execution. These include being presented with server errors such as 404 Not Found, and 403 Forbidden, encountering a captcha, and being blocked, even if just temporary, from the target website. These filtering steps produce an automatically curated set of $7,463$ demonstrations from our pipeline where agents successfully completed tasks generated by the InSTA pipeline. We reserve 500 demonstrations from this pool for our test set, and the rest for training agents in Figure 10. The original Mind2Web dataset contains $2,350$ tasks.
Appendix I Additional Related Works
While writing this paper, concurrent work was released that introduces a Proposer-Agent-Evaluator framework for web navigation agents (Zhou et al., 2024c). There are several key differences between our work and theirs, and the most important difference is scale. We generate tasks for 1M sites on the internet, whereas their work considers just 85 real-world sites, 5 sites from WebArena (Zhou et al., 2024b), and 13 sites from WebVoyager (He et al., 2024). The second difference is evaluation. Safety and reliability play crucial roles when gathering data, and we conduct an analysis on the safety and reliability of data generated by our method on 100 real-world sites. Another major difference pertains to offline learning. Offline learning should be used when scaling agents because current agent capabilities are low, and training them online risks polluting the internet with noisy LLM outputs, while taking bandwidth away from real users. The final difference pertains to the train-test split. We train agents on diverse internet data, and transfer to target benchmarks, while the agents presented in Zhou et al. (2024c) train on sites from target benchmarks using synthetic tasks. Our train-test split is stronger, and evaluates the ability for agents trained on our synthetic data to generalize to novel websites, domains, and tasks.
Appendix J Hyperparameters
We provide a list of the hyperparameters used in this work in Table 1. Values are selected to mirror prior work in synthetic data (Trabucco et al., 2024), and to employ standard hyperparameters for training agents on WebLINX (Lù et al., 2024), and Mind2Web (Deng et al., 2023). On the Mind2Web benchmark, we did not observe a loss in performance for training longer on few-shot experiments, and kept training iterations fixed between few-shot and all-available-data experiments. This is likely due to the FLAN-T5 model used in Mind2Web being too small to overfit to the data. For experiments on WebLINX, we found that performance for models trained on just human data degrades if models are trained for the full $10,000$ steps, so a smaller $2,000$ steps was chosen.
@@ Appendix I Additional Related Works
OpenAI Model Name gpt-4o
Google Model Name gemini-1.5-pro
Llama Model Name meta-llama/Llama-3.1-70B-Instruct
CommonCrawl PageRank Revision cc-main-2024-apr-may-jun-host-ranks.txt.gz
@@
Max Tokens Per Observation $4,096$
Max Tokens Per Action $2,048$
Max Tokens Per Judgement $2,048$
Max Tokens Per Task $64$
Max Observations Per Agent Context 5
Max Actions Per Agent Context 5
Max Observations Per Judge Context 1
Max Actions Per Judge Context 5
OpenAI Inference API Sampling Temperature 0.5
OpenAI Inference API Sampling Top P 1.0
Mind2Web HuggingFace Model Name google/flan-t5-large
Mind2Web Few-Shot Training Iterations $11,505$
@@
WebLINX HuggingFace Model Name meta-llama/Llama-3.1-8B-Instruct
WebLINX Few-Shot Training Iterations $2,000$
WebLINX All Training Iterations $10,000$
@@
Filtering Success Threshold 1.0
Filtering On Right Track Threshold 1.0
Few-Shot $p_{\text{real}}$ 50%
All Human Data $p_{\text{real}}$ 80%
Table 1: Hyperparameters used in our paper. We organize hyperparameters into five sections, including names of language model backbones, parameters of the data generation pipeline, sampling parameters for the OpenAI inference API, training parameters used by corresponding benchmarks, and filtering parameters used to prepare our data for training agents in Section 6.
Appendix K Cost Analysis For Llama 3.1 70B
To contextualize how Llama 3.1 70B is important for a project at this scale, we analyze the number of tokens processed by the LLM, and compute an expected cost if this were served using proprietary models. As the analysis shows, using Llama 3.1 70B is most feasible option for running agents at this large scale, and results in the paper show that this choice of LLM backbone does not compromise in accuracy and performance. We have deep gratitude for the Llama team at Meta working to make developments in language modeling available to the research community at no cost.
@@
Typical tokens per observation $1,024$
@@
Typical agent / judge response size $128$
@@
Max steps per task $10$
Estimated tokens processed by the agent
$146,746*(((1,024+128)*5+1,024)*10-(1,024+128)*15)=$ $7,419,477,760$
Tokens processed by the judge
$146,746*(1,024+1,024+128*10)=$ $488,370,688$
Total tokens processed $7,907,848,448$
Expected API cost for GPT-4o \$ $39,539.24$
Expected API cost for Gemini 1.5 Pro \$ $9,489.42$
@@
       (14 days for two 8-gpu v100 spot instances) \$ $6,622.56$
Percent saved using Llama 3.1 70B $[30.2,83.3]$

### v2

@@ Appendix G Agent & Judge System Prompts
We provide the system prompt that powers the agent in this paper. This prompt is released in the code, alongside a fast HTML to Markdown processor we built. The agent prompt is carefully designed to elicit reasoning capabilities, and the experiment in Figure 10 shows the prompt is successful.
You are helping me complete tasks by operating a web browser. I will share the current task, and a sequence of webpages and actions from previous steps.
## Action Instructions
Based on the information we discovered so far, and the progress we made in previous steps, you are helping me determine the next action.
You will provide an action as JSON in a fenced code block:
@@
"target_element_id": int | null
Actions have the following components:
- action_key‘: The name of the selected action.
- action_kwargs‘: A dictionary of arguments for the action.
- target_element_id‘: An optional id for the element to call the action on.
## Action Definitions
Ive prepared an API documentation below that defines the actions we can use to complete the task.
@@
Suppose you want to click ‘[id: 5] Sales link‘:
@@
Suppose you want to hover over ‘[id: 2] Company Logo image‘:
@@
### Scroll Action Definition
- scroll‘: Scroll the page by delta_x pixels to the right and delta_y pixels down.
- delta_x‘: The number of pixels to scroll to the right.
- delta_y‘: The number of pixels to scroll down.
### Example Scroll Action
Suppose you want to scroll down the page by 300 pixels:
‘‘‘json
"action_key": "scroll",
"action_kwargs": {
"delta_x": 0,
"delta_y": 300
"target_element_id": null
@@
### Example Fill Action (Text Input)
Suppose you want to fill ‘[id: 13] "Name..." (Enter your name text input)‘ with the text John Doe‘:
@@
### Example Fill Action (Range Slider)
Suppose you want to set ‘[id: 71] "\$250 (5)" (range slider min: 0 max: 50 step: 1)‘ to the value of \$1000‘. The slider has a range of 0 to 50 with a step of 1, and the value is currently set to ‘5‘. You must translate the desired \$1000 to the correct underlying value of ‘20‘:
‘‘‘json
"action_key": "fill",
"action_kwargs": {
"value": "20"
"target_element_id": 71
@@
Suppose you want to select the option red from ‘[id: 67] "blue" (color select from: red, blue, green)‘:
@@
### Set Checked Action Definition
- set_checked‘: Check or uncheck a checkbox specified by target_element_id‘.
- checked‘: Boolean value to check or uncheck the checkbox.
### Example Set Checked Action
Suppose you want to check ‘[id: 21] "I agree to the terms and conditions" (checkbox)‘:
‘‘‘json
"action_key": "set_checked",
"action_kwargs": {
"checked": true
"target_element_id": 21
@@
Suppose you want to open the DuckDuckGo search engine:
@@
- stop‘: Stop when the task is complete, and report your progress.
- answer‘: Optional answer sent back to me.
@@
Suppose the task is complete, and you want to stop and report your progress:
@@
"answer": "The desired task is now complete."
@@
## Formatting Your Response
Write a 200 word revised plan based on new information we discovered, and progress we made in previous steps. After your response, provide the next action as JSON in a fenced code block.
We also provide the system prompt used by the judge. The system prompt instructs the judge to predict JSON within a fenced code block that contains a “success” key, an “efficiency” key, and a “self_correction” key. The success key represents a score from 0 to 1 that estimates the probability the task is successfully completed. The efficiency key represents a score from 0 to 1 that estimates the probability the agent has taken the most efficient path to solve the task. The self correction key represents a score from 0 to 1 that estimates the probability that the agent has demonstrated self corrective behaviors during its completion of the task. These behaviors include when the agent backtracks to a more promising state, re-plans when new information is discovered relevant to the task, and recognizes its own mistakes. These are generally behaviors we expect from successful agents, but for this paper we only filter by the success key to select training data for agents.
You are helping me evaluate a browser automation script. I will share a task provided to the script, and a sequence of webpages and actions produced by the script.
## The Action Format
The script produces actions as JSON in a fenced code block:
@@
Actions have the following components:
- action_key‘: The name of the selected action.
- action_kwargs‘: A dictionary of arguments for the action.
- target_element_id‘: An optional id for the element to call the action on.
## Action Definitions
Ive prepared an API documentation below that defines the actions the script can use to complete the task.
### Click Action Definition
- click‘: Click on an element specified by target_element_id‘.
@@
Here is an example where the script clicks ‘[id: 5] Sales link‘:
@@
### Hover Action Definition
- hover‘: Hover over an element specified by target_element_id
### Example Hover Action
Here is an example where the script hovers over ‘[id: 2] Company Logo image‘:
‘‘‘json
"action_key": "hover",
"action_kwargs": {},
"target_element_id": 2
### Scroll Action Definition
- scroll‘: Scroll the page by delta_x pixels to the right and delta_y pixels down.
- delta_x‘: The number of pixels to scroll to the right.
- delta_y‘: The number of pixels to scroll down.
### Example Scroll Action
Here is an example where the script scrolls down the page by 300 pixels:
‘‘‘json
"action_key": "scroll",
"action_kwargs": {
"delta_x": 0,
"delta_y": 300
"target_element_id": null
### Fill Action Definition
- value‘: The text value to fill into the element.
### Example Fill Action (Text Input)
Here is an example where the script fills ‘[id: 13] "Name..." (Enter your name text input)‘ with the text John Doe‘:
‘‘‘json
"action_key": "fill",
"action_kwargs": {
"value": "John Doe"
"target_element_id": 13
### Example Fill Action (Range Slider)
Here is an example where the script sets ‘[id: 71] "\$250 (5)" (range slider min: 0 max: 50 step: 1)‘ to the value of \$1000‘. This slider has a range of 0 to 50 with a step of 1, and the value is currently set to ‘5‘. The script translates the desired \$1000 to the correct underlying value of ‘20‘:
‘‘‘json
"action_key": "fill",
"action_kwargs": {
"value": "20"
"target_element_id": 71
### Select Action Definition
- label‘: The option name to select in the element.
@@
Here is an example where the script selects the option red from ‘[id: 67] "blue" (color select from: red, blue, green)‘:
@@
### Set Checked Action Definition
- set_checked‘: Check or uncheck a checkbox specified by target_element_id‘.
- checked‘: Boolean value to check or uncheck the checkbox.
### Example Set Checked Action
Here is an example where the script checks ‘[id: 21] "I agree to the terms and conditions" (checkbox)‘:
‘‘‘json
"action_key": "set_checked",
"action_kwargs": {
"checked": true
"target_element_id": 21
### Go Back Action Definition
### Example Go Back Action
Here is an example where the script goes back to the previous page:
‘‘‘json
"action_key": "go_back",
"action_kwargs": {},
"target_element_id": null
### Goto Action Definition
- goto‘: Navigate to a new page (‘target_element_id must be null).
- url‘: The URL of the page to navigate to.
@@
Here is an example where the script opens DuckDuckGo search:
@@
### Stop Action Definition
- stop‘: Stop when the task is complete, and report the progress.
- answer‘: Optional answer from the script.
@@
Here is an example where the script stops and reports its progress:
@@
"answer": "The desired task is now complete."
@@
## Evaluation Instructions
Based on the progress of the script, you are helping me determine if the desired task has been completed successfully.
You will provide scores as JSON in a fenced code block:
@@
"efficiency": float,
"self_correction": float
### Score Definitions
- success‘: Your confidence the desired task has been completed successfully.
- range: 0.0 (not possible) to 1.0 (absolutely certain).
- efficiency‘: Your confidence the script has taken the most efficient path to complete the task.
- range: 0.0 (not possible) to 1.0 (absolutely certain).
- self_correction‘: Your confidence the script has demonstrated self-corrective behaviors during its completion of the task. These behaviors include backtracking to a more promising state, replanning when new information is discovered, and recognizing its own mistakes.
- range: 0.0 (not possible) to 1.0 (absolutely certain).
Write a 300 word analysis that establishes specific criteria to rigorously evaluate whether the task was completed, followed by which criteria the script has satisfied. After your response, provide your scores as JSON in a fenced code block.
@@ Appendix H Details For Training Agents
To understand the utility of data we obtained, we train agents and test on four relevant benchmarks: InSTA, WebVoyager (He et al., 2024), Mind2Web (Deng et al., 2023), WebLINX (Lù et al., 2024). In particular, our test set consists of a held-out set of 3,000 websites and tasks produced by the task generation feedback loop. Note these websites are not present in the training set. For WebVoyager (He et al., 2024), we transfer agents trained on our data zero-shot to 643 tasks on 15 websites WebVoyager (He et al., 2024). The websites in the WebVoyager benchmark are not present in the 20k trajectories we collected in Section 6.1 we used for training agents. For this experiment, we fine-tuned models based on Qwen 3 1.7B with a maximum sequence length of 16,384 tokens, and the most recent 5 observations, and actions in the context. We employed full fine-tuning on this model, with Adam, learning rate of 5e-5, batch size of 32, bfloat16, and other parameters kept as the PyTorch defaults for Adam. Each model was trained using one epoch, a linear warm-up corresponding to the first 0.01 steps of training, and a linear decay to 6e-5 afterwards.
To filter data, we select trajectories that were scores as Judge(Success) = 1, which corresponded to 10.5k of 20k trajectories produced by the Qwen 3 235B data collection policy. Scores for filtering were produced by a Qwen 3 235B judge. We employed a simple filtering strategy that only considers the success score from the judge, and no other filtering conditions were used. Note the judge also predicts efficiency, and self correction scores, which could likely also help select the best data for training, but we did not explore filtering by these scores in this work.
@@
For experiments on static benchmarks, we fine-tune google/flan-t5-large for Mind2Web, and meta-llama/Llama-3.1-8B-Instruct for WebLINX using official fine-tuning code released with corresponding benchmarks. We employ identical training hyperparameters to those used by Lù et al. (2024) for Llama in their official training code and Deng et al. (2023) for Flan to ensure that our results are comparable to previous work. Section 6.2 reports performance on the official test_web split of the WebLINX benchmark, and the official test_website split of the Mind2Web benchmark, where agents are tested on previously unobserved websites. The websites in these static benchmarks were not present in the dataset produced by InSTA for this experiment, to ensure fairness.
Appendix I Hyperparameters
We provide a list of the hyperparameters used in this work in Table 3. Hyperparameters for mixing our data with human data on static benchmarks are selected to mirror prior work in synthetic data (Trabucco et al., 2024), and to adhere to standard hyperparameters for WebLINX (Lù et al., 2024), and Mind2Web (Deng et al., 2023). We train using all available human data on these benchmarks, and add trajectories filtered using the previously discussed methodology, sampled at a 20% rate in the data-loader compared to an 80% rate for human data.
@@ Appendix I Hyperparameters
Models Used For Agents Qwen/Qwen3-1.7B
Qwen/Qwen3-235B-A22B
meta-llama/Llama-4-Maverick-17B-128E-Instruct
meta-llama/Llama-3.1-70B-Instruct
meta-llama/Llama-3.3-70B-Instruct
google/gemini-2.5-flash
Models Used For Judges Qwen/Qwen3-235B-A22B
meta-llama/Llama-4-Maverick-17B-128E-Instruct
meta-llama/Llama-3.1-70B-Instruct
meta-llama/Llama-3.3-70B-Instruct
google/gemini-2.5-flash
google/gemini-1.5-pro
openai/gpt-4.1-mini
openai/gpt-4o
Common Crawl PageRank cc-main-2024-apr-may-jun-host-ranks.txt.gz
@@
Max Tokens Per Observation $2,048$
Max Tokens Per Agent Trace $1,024$
Max Tokens Per Judge Trace $1,024$
Max Tokens Per Task Proposer Trace $1,024$
Last Steps Per Agent Context 5
Last Steps Per Judge Context 5
Last Steps Per Task Proposer Context 5
Task Proposer Feedback Loops 1
LLM Sampling Temperature 0.5
LLM Sampling Top P 1.0
LLM Sampling Top K default
Fine-tuned LLM in Section 6.1 Qwen/Qwen3-1.7B
InSTA Training Epochs 1
InSTA Batch Size $32$
InSTA Learning Rate 5e-5
InSTA Optimizer Adam
Mind2Web LLM google/flan-t5-large
@@
Mind2Web Optimizer Adam
WebLINX LLM meta-llama/Llama-3.1-8B-Instruct
WebLINX Training Iterations $10,000$
@@
WebLINX Optimizer Adam
Data Filtering Condition Judge(Success) = 1
Human Data Sampling Probability $p_{\text{real}}$ 80%
Table 3: Hyperparameters used in our paper. We organize hyperparameters into seven sections, for the names of LLMs used as agents in the paper, the names of LLMs used as judges in the paper, the hyperparameters used for data collection, the sampling parameters for LLMs, the training parameters for static benchmarks, and the filtering and data mixing hyperparameters.
Appendix J Cost Analysis For Llama 3.1 70B
To understand the significant reduction in cost that we obtain by running LLMs locally to generate data, we analyze the number of tokens processed by the LLM, and compute an expected cost if this were served using proprietary models. As the analysis shows, using Llama 3.1 70B is a feasible option for running agents at this large scale, and results in the paper show that this choice of LLM backbone does not compromise performance. We have deep gratitude to the Llama team at Meta, and the Qwen team at Alibaba for working to make developments in language modeling available to the research community at no cost. We see up to a 95% reduction in cost with these models.
@@
Average tokens per observation $1,024$
@@
Average agent / judge response size $512$
@@
Average steps per task $15$
Estimated tokens processed by the agent 14.65B tokens
Tokens processed by the judge 1.35B tokens
Total tokens processed 16.00B tokens
Expected API cost for GPT-4.1 \$ 32,000.00
Expected API cost for Gemini 2.5 Pro \$ 20,000.00
@@
       (3,840 v100 GPU hours using spot instances) \$ $1,575.70$
Percent saved using Llama 3.1 70B $[95.08,92.12]$